# Differentially Private Bilevel Optimization: Efficient Algorithms with Near-Optimal Rates

### v1

@@ Abstract
Bilevel optimization, in which one optimization problem is nested inside another, underlies many machine learning applications with a hierarchical structure—such as meta-learning and hyperparameter optimization. Such applications often involve sensitive training data, raising pressing concerns about individual privacy. Motivated by this, we study differentially private bilevel optimization. We first focus on settings where the outer-level objective is convex, and provide novel upper and lower bounds on the excess risk for both pure and approximate differential privacy, covering both empirical and population-level loss. These bounds are nearly tight and essentially match the optimal rates for standard single-level differentially private ERM and stochastic convex optimization (SCO), up to additional terms that capture the intrinsic complexity of the nested bilevel structure. The bounds are achieved in polynomial time via efficient implementations of the exponential and regularized exponential mechanisms. A key technical contribution is a new method and analysis of log-concave sampling under inexact function evaluations, which may be of independent interest. In the non-convex setting, we develop novel algorithms with state-of-the-art rates for privately finding approximate stationary points. Notably, our bounds do not depend on the dimension of the inner problem.
@@ 1 Introduction
[…]
@@
[…]
@@
Question 1. What are the optimal error rates for solving problems Bilevel ERM and Bilevel SO with DP when $\widehat{\Phi}_{Z}$ and $\Phi$ are convex?
@@
We give a (nearly) complete answer to Question 1 for both pure $\varepsilon$-DP and approximate $(\varepsilon,\delta)$-DP, by providing nearly tight upper and lower bounds : see Section 3. Our results show that if the smoothness, Lipschitz, and strong convexity parameters are constants, then it is possible to achieve the same rates as standard single-level convex DP-ERM bst14 and DP-SO bft19 , despite the more challenging bilevel setting (e.g., $O(d_{x}/\varepsilon n)$ for $\varepsilon$-DP bilevel ERM). On the other hand, our lower bound establishes a novel separation between standard single-level DP optimization and DP BLO, showing that the error of any algorithm for DP BLO must necessarily depend on the complexity parameters of the lower-level problem (e.g. the Lipschitz parameter of $g(x,\cdot,z)$). Our algorithms are built on the exponential mechanism mcsherry2007mechanism for $\varepsilon$-DP and the regularized exponential mechanism gopi2022private for $(\varepsilon,\delta)$-DP. We provide efficient (i.e. polynomial-time) implementations of these mechanisms for DP BLO and a novel analysis of how function evaluation errors affect log-concave sampling algorithms.
@@ 1.1 Technical overview
In the nonconvex setting, our algorithm uses a second-order approximation ¯$\bar{\nabla}\widehat{F}_{Z}(x_{t},y_{t+1})\approx\nabla\widehat{\Phi}_{Z}(x_{t})$ in order to approximate gradient descent run on $\widehat{\Phi}_{Z}$. A key insight is that we can obtain a better bound by getting a high-accuracy non-private approximate solution $y_{t+1}\approx\widehat{y}^{*}_{Z}(x_{t})$ and then noising ¯$\bar{\nabla}\widehat{F}_{Z}(x_{t},y_{t+1})$, rather than privatizing $y_{t+1}$. To prove such an approach can be made DP, we require a careful sensitivity analysis that leverages perturbation inequalities from numerical analysis. Further, we build a two-step algorithm on our novel second-order algorithm by leveraging the warm-start framework of lowy2024make .
@@ 2 Preliminaries
[…]
@@
Note that $\text{diam}(\mathcal{Y})\leq\frac{L_{g,y}}{\mu_{g}}$ by Assumption 2.1. Some of our algorithms additionally require:
@@ 3 Private convex bilevel optimization
In this section, we characterize the optimal excess risk bounds for DP convex bilevel ERM and SCO:
@@
$\mathbb{E}\widehat{\Phi}_{Z}(\widehat{x})-\widehat{\Phi}_{Z}^{*}\leq O\left(\frac{\sqrt{d_{x}\log(1/\delta)}}{\varepsilon n}\right),~{}~{}\text{and}$
$\mathbb{E}\Phi(\widehat{x})-\Phi^{*}\leq O\left(\frac{\sqrt{d_{x}\log(1/\delta)}}{\varepsilon n}+\frac{1}{\sqrt{n}}\right).$
Moreover, all of the above upper bounds are tight (optimal) up to logarithmic factors.
The following subsections contain formal statements capturing the precise dependence on the problem parameters given in Assumptions 2.1 and 2.2 and runtime bounds.
@@ 3.1 Conceptual algorithms and excess risk upper bounds
$\mathbb{E}\Phi(\widehat{x})-\Phi^{*}\leq O\left(\left(L_{f,x}+\frac{L_{f,y}\beta_{g,xy}}{\mu_{g}}+\frac{L_{g,y}\beta_{f,xy}}{\mu_{g}}\right)D_{x}\left(\frac{\sqrt{d_{x}\log(1/\delta)}}{\varepsilon n}+\frac{1}{\sqrt{n}}\right)\right).$
The main idea of the privacy proof (in Appendix B) is to show that $\widehat{\Phi}_{Z}-\widehat{\Phi}_{Z^{\prime}}$ is $2(\frac{L_{f,x}}{n}+\frac{L_{f,y}\beta_{g,xy}}{\mu_{g}n}+\frac{L_{g,y}\beta_{f,xy}}{n\mu_{g}})$-Lipschitz and then compare the privacy curve balle2018improving between the distributions $Q$ and $Q^{\prime}$ (corresponding to (5) with data $Z$ and $Z^{\prime}$ respectively) to the privacy curve between two Gaussians, by leveraging (gopi2022private, , Theorem 4.1).
Remark 3.4 (Near-optimality).
The bounds in Theorem 3.2 and 3.3 nearly match the optimal bounds for standard single-level DP ERM and SCO bst14 ; bft19 , e.g. $\Theta(L_{f,x}D_{x}\sqrt{d_{x}\log(1/\delta)}/\varepsilon n)$ for $(\varepsilon,\delta)$-DP ERM bst14 ; su16 , except for the addition of two terms capturing the complexity of the bilevel problem: For $\varepsilon$-DP ERM, the additional terms are $O(L_{f,y}D_{y}d_{x}/\varepsilon n)$ and $O((L_{f,y}L_{g,y}/\mu_{g})d_{x}/\varepsilon n)$. Our lower bound in Theorem 3.9 shows that the first additional term is necessary. We conjecture that the second additional term is also necessary and that our upper bound is tight up to an absolute constant. This conjecture is clearly true in the parameter regime $L_{g,y}/\mu_{g}\approx D_{y}$. For $(\varepsilon,\delta)$-DP, the additional terms scale with $O((L_{f,y}\beta_{g,xy}/\mu_{g}+L_{g,y}\beta_{f,xy}/\mu_{g})D_{x})$. Our lower bound in Theorem 3.9 shows that dependence on $L_{f,y}$ is necessary and that the $L_{f,y}\beta_{g,xy}/\mu_{g}$ term is tight in the parameter regime $D_{y}\approx D_{x}\beta_{g,xy}/\mu_{g}$. If also $D_{x}\beta_{g,xy}/\lesssim L_{g,y}$, then the bounds in Theorem 3.3 are tight up to an absolute constant factor.
@@ 3.3 Excess risk lower bounds
One can obtain lower bounds on the excess population risk that are larger than the excess empirical risk bounds in Theorem 3.9 by an additive $L_{f,x}D_{x}(1/\sqrt{n})$, via the reduction in BassilyFeTaTh19 . This implies the near-optimality of our scheme in (5) for DP bilevel convex SO (c.f. Theorem 3.3).
@@ 4.1 An iterative second-order method
¯$\bar{\nabla}\widehat{F}_{Z}(x,y):=\nabla_{x}\widehat{F}_{Z}(x,y)-\nabla_{xy}^{2}\widehat{G}_{Z}(x,y)[\nabla_{yy}^{2}\widehat{G}_{Z}(x,y)]^{-1}\nabla_{y}\widehat{F}_{Z}(x,y).$ (9)
Note that ¯$\bar{\nabla}\widehat{F}_{Z}(x,y)=\nabla\widehat{\Phi}_{Z}(x)$ if $y=\widehat{y}^{*}_{Z}(x)$.
@@
A naive approach to privatizing the iterations (4.1) is to solve $y_{t+1}\approx\widehat{y}^{*}_{Z}(x_{t})=\text{argmin}_{y}\widehat{G}_{Z}(x_{t},y)$ privately at each step (e.g., by running DP-SGD), and then add noise to ¯$\bar{\nabla}\widehat{F}_{Z}(x_{t},y_{t+1})$ before taking a step of noisy GD. (This is similar to how kornowski2024differentially privatized the penalty-based bilevel optimization algorithm of kwon2023fully .) However, this approach results in a bound $\mathbb{E}\|\nabla\widehat{\Phi}_{Z}(\widehat{x})\|\leq O(\sqrt{d_{x}+d_{y}}/\varepsilon n)^{1/2}$ that depends on $d_{y}$ due to the bias ¯$\|\bar{\nabla}\widehat{F}_{Z}(x_{t},y_{t+1})-\nabla\widehat{\Phi}_{Z}(x_{t})\|$ that results from using private $y_{t+1}$. To mitigate this issue and obtain state-of-the-art utility independent of $d_{y}$, we propose an alternative approach in Algorithm 1: we find an approximate minimizer of $\widehat{G}_{Z}(x_{t},\cdot)$ non-privately in line 3. Since $\widehat{G}_{Z}(x_{t},\cdot)$ is a smooth, strongly convex ERM function, we can implement line 3 efficiently using a non-private algorithm such as SGD or Katyusha allen2017katyusha .
1 Input: Dataset $\mathcal{D}=(Z_{1},\dots,Z_{n})$, noise scale $\sigma$, initial points $x_{0},y_{0}\in\mathcal{X}\times\mathcal{Y}$, parameter $\alpha$;
@@
The proof of this lemma—in Appendix C—is long. It uses the operator norm perturbation inequality $\|M^{-1}-N^{-1}\|\leq\|M^{-1}\|\|N^{-1}\|\|M-N\|$ to bound the sensitivity of $[\nabla^{2}_{yy}\widehat{G}_{Z}(x_{t},y_{y+1})]^{-1}$ in (9).
@@
Grant Assumptions 2.1 and 2.2. Set $\sigma=32K\sqrt{T\log(1/\delta)}/n\varepsilon$. Denote the smoothness parameter of $\widehat{\Phi}_{Z}$ by $\beta_{\Phi}$, given in Lemma C.2. There are choices of $\alpha,\eta$ s.t. Algorithm 1 is $(\varepsilon,\delta)$-DP and has output satisfying
@@
Grant Assumptions 2.1 and 2.2. Assume that there is a compact set $\mathcal{X}\subset\mathbb{R}^{d_{x}}$ of diameter $D_{x}$ containing an approximate global minimizer $\widehat{x}$ such that $\widehat{\Phi}_{Z}(\widehat{x})-\widehat{\Phi}_{Z}^{*}\leq\Psi\frac{d}{\varepsilon n}$, where $\Psi:=L_{f,x}D_{x}+L_{f,y}D_{y}+\frac{L_{f,y}L_{g,y}}{\mu_{g}}.$ Then, there exists an $(\varepsilon,\delta)$-DP instantiation of Algorithm 2 with output satisfying

### v2

@@ Abstract
Bilevel optimization, in which one optimization problem is nested inside another, underlies many machine learning applications with a hierarchical structure—such as meta-learning and hyperparameter optimization. Such applications often involve sensitive training data, raising pressing concerns about individual privacy. Motivated by this, we study differentially private bilevel optimization. We first focus on settings where the outer-level objective is convex, and provide novel upper and lower bounds on the excess empirical risk for both pure and approximate differential privacy. These bounds are nearly tight and essentially match the optimal rates for standard single-level differentially private ERM, up to additional terms that capture the intrinsic complexity of the nested bilevel structure. We also provide population loss bounds for bilevel stochastic optimization. The bounds are achieved in polynomial time via efficient implementations of the exponential and regularized exponential mechanisms. A key technical contribution is a new method and analysis of log-concave sampling under inexact function evaluations, which may be of independent interest. In the non-convex setting, we develop novel algorithms with state-of-the-art rates for privately finding approximate stationary points. Notably, our bounds do not depend on the dimension of the inner problem.
@@ 1 Introduction
$\text{s.t.}~~\widehat{y}^{*}_{Z}(x)=\text{argmin}_{y\in\mathbb{R}^{d_{y}}}\left\{\widehat{G}_{Z}(x,y)=\frac{1}{n}\sum_{i=1}^{n}g(x,\widehat{y}^{*}_{Z}(x),z_{i})\right\},$
@@
$\text{s.t.}~~y^{*}(x)=\text{argmin}_{y\in\mathbb{R}^{d_{y}}}\left\{G(x,y)=\mathbb{E}_{z\sim P}[g(x,y,z)]\right\}.$
@@
Question 1. What are the optimal error rates for solving problem Bilevel ERM with DP when $\widehat{\Phi}_{Z}$ is convex?
Convex $\widehat{\Phi}_{Z},\Phi$ arise in a variety of applications liang2023lower , including few-shot meta-learning with a shared embedding model bertinetto2018meta , biased regularization in hyperparameter optimization grazzi2020iteration , fair resource allocation in communication networks srikant2014communication , and bilevel optimization with smooth convex $f(\cdot,y)$ and quadratic $g(x,\cdot)$ liang2023lower .
@@
We give a (nearly) complete answer to Question 1 for both pure $\varepsilon$-DP and approximate $(\varepsilon,\delta)$-DP, by providing nearly tight upper and lower bounds: see Section 3. Our results show that if the smoothness, Lipschitz, and strong convexity parameters are constants, then it is possible to achieve the same rates as standard single-level convex DP-ERM bst14 , despite the more challenging bilevel setting (e.g., $O(d_{x}/\varepsilon n)$ for $\varepsilon$-DP bilevel ERM). On the other hand, our lower bound establishes a novel separation between standard single-level DP optimization and DP BLO, showing that the error of any algorithm for DP BLO must necessarily depend on the complexity parameters of the lower-level problem (e.g. the Lipschitz parameter of $g(x,\cdot,z)$). Our algorithms are built on the exponential mechanism mcsherry2007mechanism for $\varepsilon$-DP and the regularized exponential mechanism gopi2022private for $(\varepsilon,\delta)$-DP. We provide efficient (i.e. polynomial-time) implementations of these mechanisms for DP BLO and a novel analysis of how function evaluation errors affect log-concave sampling algorithms. Additionally, we provide upper and lower bounds on the excess population risk for DP Bilevel SO.
@@ 1.1 Technical overview
In the nonconvex setting, our algorithm uses a second-order approximation $\bar{\nabla}\widehat{F}_{Z}(x_{t},y_{t+1})\approx\nabla\widehat{\Phi}_{Z}(x_{t})$ in order to approximate gradient descent run on $\widehat{\Phi}_{Z}$. A key insight is that we can obtain a better bound by getting a high-accuracy non-private approximate solution $y_{t+1}\approx\widehat{y}^{*}_{Z}(x_{t})$ and then noising $\bar{\nabla}\widehat{F}_{Z}(x_{t},y_{t+1})$, rather than privatizing $y_{t+1}$. To prove such an approach can be made DP, we require a careful sensitivity analysis that leverages perturbation inequalities from numerical analysis. Further, we build a two-step algorithm on our novel second-order algorithm by leveraging the warm-start framework of lowy2024make .
@@ 2 Preliminaries
Let $f:\mathcal{X}\times\mathbb{R}^{d_{y}}\times\mathcal{Z}\to\mathbb{R}$ and $g:\mathcal{X}\times\mathbb{R}^{d_{y}}\times\mathcal{Z}\to\mathbb{R}$ be loss functions, with $\mathcal{X}\subset\mathbb{R}^{d_{x}}$ being a closed convex set of $\ell_{2}$-diameter $D_{x}\in[0,\infty]$. The data universe $\mathcal{Z}$ can be any set. $P$ denotes any data distribution on $\mathcal{X}$. Let $\|\cdot\|$ denote the $\ell_{2}$ norm when applied to vectors. When applied to matrix $A$, $\|A\|:=s_{\max}(A)=\sqrt{\lambda_{\max}(AA^{T})}$ denotes the $\ell_{2}$ operator norm, which is the largest singular value of $A$. Function $h:\mathcal{X}\to\mathbb{R}$ is $L$-Lipschitz if $|h(x)-h(x^{\prime})|\leq L\|x-x^{\prime}\|$ for all $x,x^{\prime}\in\mathcal{X}$. Function $h:\mathcal{X}\to\mathbb{R}$ is $\mu$-strongly convex if $h(\alpha x+(1-\alpha)x^{\prime})\leq\alpha h(x)+(1-\alpha)h(x^{\prime})-\frac{\alpha(1-\alpha)\mu}{2}\|x-x^{\prime}\|^{2}$ for all $\alpha\in[0,1]$ and all $x,x^{\prime}\in\mathcal{X}$. If $\mu=0,$ we say $h$ is convex. The excess (population) risk of a randomized algorithm $\mathcal{A}$ with output $\widehat{x}=\mathcal{A}(Z)$ on loss function $h(x,z)$ is $\mathbb{E}_{\mathcal{A},Z}[H(\widehat{x})]-H^{*}$, where $H(x)=\mathbb{E}_{z\sim P}[h(x,z)]$ and $H^{*}:=\inf_{x}H(x)$. If $\widehat{H}_{Z}(x)=\frac{1}{n}\sum_{i=1}^{n}h(x,z_{i})$ is an empirical loss function w.r.t. data set $Z$, then the excess empirical risk of $\mathcal{A}$ is $\mathbb{E}_{\mathcal{A}}[\widehat{H}_{Z}(\widehat{x})]-\widehat{H}^{*}$. Denote $a\wedge b:=\min(a,b)$. For functions $\varphi$ and $\psi$ of input parameters $\theta$, we write $\varphi\lesssim\psi$ if there is an absolute constant $C>0$ such that $\varphi(\theta)\leq C\psi(\theta)$ for all permissible values of $\theta$. We use $\widetilde{O}$ to hide logarithmic factors. Denote by $\nabla J(x,y(x),z)=\nabla_{x}J(x,y(x),z)+\nabla y(x)^{T}\nabla_{y}J(x,y(x),z)$ the gradient of function $J$ w.r.t. $x$.
@@
Note that $D_{y}:=\text{diam}(\mathcal{Y})\leq\frac{L_{g,y}}{\mu_{g}}$ by Assumption 2.1. Some of our algorithms additionally require:
@@ 3 Private convex bilevel optimization
In this section, we characterize the optimal excess risk bounds for DP convex bilevel ERM:
@@
$\mathbb{E}\widehat{\Phi}_{Z}(\widehat{x})-\widehat{\Phi}_{Z}^{*}\leq O\left(\frac{\sqrt{d_{x}\log(1/\delta)}}{\varepsilon n}\right)$
Moreover, the above upper bounds are tight (optimal) up to logarithmic factors.
The following subsections contain formal statements capturing the precise dependence on the problem parameters given in Assumptions 2.1 and 2.2 and runtime bounds. We also provide bounds for DP convex bilevel SO.
@@ 3.1 Conceptual algorithms and excess risk upper bounds
$\mathbb{E}\Phi(\widehat{x})-\Phi^{*}\leq O\left(\frac{\sqrt{d_{x}\log(1/\delta)}}{\varepsilon n}+\frac{1}{n^{1/4}}\right).$
Refer to Theorem B.8 in Appendix B for the precise population loss bound with proper units. The main idea of the privacy proof (in Appendix B) is to show that $\widehat{\Phi}_{Z}-\widehat{\Phi}_{Z^{\prime}}$ is $2(\frac{L_{f,x}}{n}+\frac{L_{f,y}\beta_{g,xy}}{\mu_{g}n}+\frac{L_{g,y}\beta_{f,xy}}{n\mu_{g}})$-Lipschitz and then compare the privacy curve balle2018improving between the distributions $Q$ and $Q^{\prime}$ (corresponding to (5) with data $Z$ and $Z^{\prime}$ respectively) to the privacy curve between two Gaussians, by leveraging (gopi2022private, , Theorem 4.1).
Obtaining our dimension-independent generalization error bound involves a fairly long “ghost sample” argument that leverages the Wasserstein distance bound for log-concave distributions, Kantorovich-Rubinstein duality, and the Efron-Stein inequality. One can also obtain an $O(\sqrt{d/n})$ generalization error bound by a uniform convergence argument; we omit those details here.
Remark 3.4 (Near-optimality for ERM).
The bounds in Theorem˜3.2 and 3.3 nearly match the optimal bounds for standard single-level DP ERM bst14 , e.g. $\Theta(L_{f,x}D_{x}\sqrt{d_{x}\log(1/\delta)}/\varepsilon n)$ for $(\varepsilon,\delta)$-DP ERM bst14 ; su16 , except for the addition of two terms capturing the complexity of the bilevel problem: For $\varepsilon$-DP ERM, the additional terms are $O(L_{f,y}D_{y}d_{x}/\varepsilon n)$ and $O((L_{f,y}L_{g,y}/\mu_{g})d_{x}/\varepsilon n)$. Our lower bound in Theorem 3.10 shows that the first additional term is necessary. We conjecture that the second additional term is also necessary and that our upper bound is tight up to an absolute constant. This conjecture is clearly true in the parameter regime $L_{g,y}/\mu_{g}\approx D_{y}$. For $(\varepsilon,\delta)$-DP, the additional terms scale with $O((L_{f,y}\beta_{g,xy}/\mu_{g}+L_{g,y}\beta_{f,xy}/\mu_{g})D_{x})$. Our lower bound in Theorem 3.10 shows that dependence on $L_{f,y}$ is necessary and that the $L_{f,y}\beta_{g,xy}/\mu_{g}$ term is tight in the parameter regime $D_{y}\approx D_{x}\beta_{g,xy}/\mu_{g}$. If also $D_{x}\beta_{g,xy}/\lesssim L_{g,y}$, then the bounds in Theorem˜3.3 are tight up to an absolute constant factor.
Remark 3.5 (Suboptimality for SO).
There is a gap between our population risk upper bound in Theorem˜3.3 and the lower bounds in bft19 and Remark 3.11 for single-level SCO and bilevel SO respectively. We conjecture that our lower bound in Remark 3.11 is nearly tight and that our upper bound is suboptimal. We leave it as future work to investigate this conjecture.
@@ 3.3 Excess risk lower bounds
One can obtain lower bounds on the excess population risk that are larger than the excess empirical risk bounds in Theorem 3.10 by an additive $L_{f,x}D_{x}(1/\sqrt{n})$, via the reduction in BassilyFeTaTh19 .
@@ 4.1 An iterative second-order method
$\bar{\nabla}\widehat{F}_{Z}(x,y):=\nabla_{x}\widehat{F}_{Z}(x,y)-\nabla_{xy}^{2}\widehat{G}_{Z}(x,y)[\nabla_{yy}^{2}\widehat{G}_{Z}(x,y)]^{-1}\nabla_{y}\widehat{F}_{Z}(x,y).$ (9)
Note that $\bar{\nabla}\widehat{F}_{Z}(x,y)=\nabla\widehat{\Phi}_{Z}(x)$ if $y=\widehat{y}^{*}_{Z}(x)$.
@@
A naive approach to privatizing the iterations (4.1) is to solve $y_{t+1}\approx\widehat{y}^{*}_{Z}(x_{t})=\text{argmin}_{y}\widehat{G}_{Z}(x_{t},y)$ privately at each step (e.g., by running DP-SGD), and then add noise to $\bar{\nabla}\widehat{F}_{Z}(x_{t},y_{t+1})$ before taking a step of noisy GD. (This is similar to how kornowski2024differentially privatized the penalty-based bilevel optimization algorithm of kwon2023fully .) However, this approach results in a bound $\mathbb{E}\|\nabla\widehat{\Phi}_{Z}(\widehat{x})\|\leq O(\sqrt{d_{x}+d_{y}}/\varepsilon n)^{1/2}$ that depends on $d_{y}$ due to the bias $\|\bar{\nabla}\widehat{F}_{Z}(x_{t},y_{t+1})-\nabla\widehat{\Phi}_{Z}(x_{t})\|$ that results from using private $y_{t+1}$. To mitigate this issue and obtain state-of-the-art utility independent of $d_{y}$, we propose an alternative approach in Algorithm 1: we find an approximate minimizer of $\widehat{G}_{Z}(x_{t},\cdot)$ non-privately in line 3. Since $\widehat{G}_{Z}(x_{t},\cdot)$ is a smooth, strongly convex ERM function, we can implement line 3 efficiently using a non-private algorithm such as SGD or Katyusha allen2017katyusha .
1Input: Dataset $\mathcal{D}=(Z_{1},\dots,Z_{n})$, noise scale $\sigma$, initial points $x_{0},y_{0}\in\mathcal{X}\times\mathcal{Y}$, parameter $\alpha$;
@@
The proof of this lemma—in Appendix C—is long. It uses the operator norm perturbation inequality $\|M^{-1}-N^{-1}\|\leq\|M^{-1}\|\|N^{-1}\|\|M-N\|$ to bound the sensitivity of $\nabla^{2}_{yy}\widehat{G}_{Z}(x_{t},y_{y+1})]^{-1}$ in (9).
@@
Grant Assumptions 2.1 and $2.$ Set $\sigma=32K\sqrt{T\log(1/\delta)}/n\varepsilon$. Denote the smoothness parameter of $\widehat{\Phi}_{Z}$ by $\beta_{\Phi}$, given in Lemma C.2. There are choices of $\alpha,\eta$ s.t. Algorithm 1 is $(\varepsilon,\delta)$-DP and has output satisfying
@@
Grant Assumptions 2.1 and $2.$ Assume that there is a compact set $\mathcal{X}\subset\mathbb{R}^{d_{x}}$ of diameter $D_{x}$ containing an approximate global minimizer $\widehat{x}$ such that $\widehat{\Phi}_{Z}(\widehat{x})-\widehat{\Phi}_{Z}^{*}\leq\Psi\frac{d}{\varepsilon n}$, where $\Psi:=L_{f,x}D_{x}+L_{f,y}D_{y}+\frac{L_{f,y}L_{g,y}}{\mu_{g}}.$ Then, there exists an $(\varepsilon,\delta)$-DP instantiation of Algorithm 2 with output satisfying